write the following sets by listing their elements between braces


1
Expert's answer
2021-09-15T03:16:46-0400

A.

2. "\{3x+2:x\in Z\}=\{...,-4,-1,2,5,8,11,14,...\}."

4.

"\{x\in N:-2<x\leq7\}=\{-1,0,1,2,3,4,5,6,7\}."

6.

"\{x \in R:x^2=9\}=\{-3,3\}."

8.

"\{x\in R:x^3+5x^2=-6x\}=\{-3,-2,0\}."

10.

"\{x\in R:\sin{πx}=0\}=\{...,-3,-2,-1,0,1,2,3,...\}."

B.

18.

"\{0,4,16,36,64,100,...\}=\{x\in Z:(2x)^2\}."

20.

"\{...,-8,-3,2,7,12,17,...\}=\{x\in Z:5x-3\}."

22.

"\{3,6,11,18,27,38,...\}=\{x\in N:x^2+2\}."

24."\{-4,-3,-2,-1,0,1,2\}=\{x \in Z:-4\leq x\leq 2\}."

26.

"\{...,\frac{1}{27},\frac{1}{9},\frac{1}{3},1,3,9,27,...\}=\{x \in Z:3^x\}."

C.

30.

"|\{\{1,4\},a,b,\{\{3,4\}\},\{\varnothing\}\}|=5."

32.

"|\{\{\{1,4\},a,b,\{\{3,4\}\},\{\varnothing\}\}\}|=1."

34.

"|\{x\in N:|x|<10\}|=9."

36.

"|\{x\in N:x^2<10\}|=3."

38.

"|\{x \in N:5x\leq 20\}|=4."